# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 4.3 Wall-crossing
which can be viewed as the tangent cone to the stratum parameterizing equigeneric deformations with the fixed intersection multiplicity $2$ with ${\operatorname{Tor}}(\sigma)$ at $w$, respectively, at a nearby point on ${\operatorname{Tor}}(\sigma)$ (see [13, Lemma 2.4(1)] and [24, Lemma 3(1)]). The required transversality is as follows: the natural image of the germ of the linear system $|{\mathcal{L}}_{P_{\Delta}}|$ at $C$ to $\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}{\mathcal{O}}_{C,z}/{\mathfrak{m}}_{z}^{s}\times\prod_{w\in{\boldsymbol{w}}(t^{*})}{\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$ intersects there transversally with $\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}I_{z}\times\prod_{w\in\widehat{\boldsymbol{w}}(t^{*})}I_{w}$. The cohomological reformulation amounts to the relation
$H^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}({\mathbb{P}}^{1})\ ,$ (12)
where the points of ${\boldsymbol{w}}(t^{*})$ are fixed and the remaining point of $\widehat{\boldsymbol{w}}(t^{*})$ is mobile. Since
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}2\delta(C,z)-\sum_{w\in\widehat{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-1)=-1>-2\ ,$
and hence (12) by the Riemann-Roch formula. The established transversality reduces the constancy of $W_{0}^{\kappa}(t)$ to the constancy of the count of Welschinger signs when varying germs $(C,z)$ and $(C,w)$ for $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$ and $w\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap\widehat{\boldsymbol{w}}(t^{*})$. Thus, the constancy in the former variation follows from [12, Lemma 15] (see also [12, Lemma 13]), while the constancy in the latter variation follows from the fact that the equigeneric stratum, which we consider in ${\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$, is smooth (see [24, Lemma 3(1)] and [13, Lemma 2.4(2)]), and the local count of Welschinger signs is invariant by [13, Lemma 2.4(1)].
(2) Suppose that $\xi\in\overline{\mathcal{M}}^{{\mathbb{R}},+}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ is as in Lemma 4.2(2i) with a singularity at a point $w_{i}^{\sigma}(t^{*})$ of type $A_{2m}$, $m\geq k_{i}^{\sigma}$. We can assume that the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<{\varepsilon}}$ is such that $n-1$ points of $\widehat{\boldsymbol{w}}(t)$, including $w_{i}^{\sigma}(t)$, stay fixed, whereas the other two points move keeping the Menelaus condition. We claim that the germ at $\xi$ of the family {¯$\{\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<{\varepsilon}}$ is smooth and regularly parameterized by $t\in(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$. By [3, Theorem 4.15 and Proposition 4.17(2)] and [24, Lemma 3(1)], it is sufficient to prove that
$H^{1}(C,{\mathcal{J}}_{Z/C}(C))=0\ ,$ (13)
where $C={\Phi}({\mathbb{P}}^{1})$ and ${\mathcal{J}}_{Z/C}\subset{\mathcal{O}}_{C}$ is the ideal sheaf of the zero-dimensional scheme $Z\subset C$ concentrated at $\widehat{\boldsymbol{w}}(t^{*})\cup{\operatorname{Sing}}\hskip 1.70709pt(C)$ and given (cf. Step (1) of the proof)
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}\}$ at each point $w=w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$, which is fixed as $|t-t^{*}|<{\varepsilon}$,
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}-1\}$ at each point $w=w_{i^{\prime}}^{\sigma^{\prime}}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$, which moves as $|t-t^{*}|<{\varepsilon}$,
  • by the ideal $I_{w}=\{\varphi\in{\mathcal{O}}_{C,w}\ :\ {\operatorname{ord}}\varphi\big{|}_{w}\geq 2k_{i}^{\sigma}+2m\}$ at the point $w=w_{i}^{\sigma}(t^{*})$ (here, in the notations of [24, Lemma 3(1)], $s=2k_{i}^{\sigma}$ and $\delta(C,w)=m$),
  • by the conductor ideal $I^{cond}_{C,z}$ at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\setminus\{w_{i}^{\sigma}(t^{*})\}$ (for the definition of the conductor ideal see, for instance, [3, Section 1, item (iii)] and [9, Section I.3.4, item “Semigroup and Conductor”]).
Thus, the desired relation (13) turns into
$H^{1}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}({\boldsymbol{d}}))=0\ ,$
(cf. (12)), where the divisor ${\boldsymbol{d}}$ has degree
$\deg{\boldsymbol{d}}=C^{2}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\setminus{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,z}/I_{z}-\sum_{w\in{\boldsymbol{w}}(t^{*})}\dim{\mathcal{O}}_{C,w}/I_{w}$
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-2)=0>-2\ ,$
which, finally, confirms (13).
[…]
$F_{t_{1}}(x,y):=F(x,y)+\sum_{j=0}^{k-1}t_{1}^{(k-j)\gamma}(a_{j1}+O(t_{1}))x^{j}y+\sum_{p+kq\geq 2k}b_{pq}(t_{1})x^{p}y^{q}=0\ ,$ (14)
where $a_{01}\neq 0$ and $b_{pq}(0)=0$ for all $p,q$ in the range. The tropical limit of the germs $[{\Phi}_{t}({\mathbb{P}}^{1}),w]$ as $t\to t^{*}$ or, equivalently, $t_{1}\to 0$ defines the subdivision of the Newton polygon $P(F_{t})$ of $F_{t}(x,y)$ into the triangle $T_{1}={\operatorname{conv}}\{(0,1),(0,2),(2k,0)\}$ and the Newton polygon $P(F)$ of $F$, see Figure 1(a) (the details about the tropicalization can be found in [18, Section 3] and [15, Section 2.5.4], see also [22, Section 2] and [5, Section 2.2]). Note that, since the curves ${\Phi}_{t}({\mathbb{P}}^{1})$ are rational, the triangle $T_{1}$ is not subdivided further, and the corresponding limit curve
$C_{1}:=\left\{\sum_{j=0}^{k-1}a_{j1}x^{j}y+y^{2}-2x^{k}y+x^{2k}=0\right\}$ (15)
is rational. Moreover, since the curve $C$ is unibranch at $w$, the limit curve $C_{F}$ defined by the polynomial $F$ in the surface ${\operatorname{Tor}}(P(F))$ is unibranch at the intersection point with the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$. Hence, $C_{1}$ is unibranch at its intersection point with the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$, because otherwise, the union of the limit curves $C_{1}$ and $C_{F}$ would deform into a non-rational curve contrary to the fact that the curves ${\Phi}({\mathbb{P}}^{1})$ are rational. By [22, Lemma 3.5], the curve $C_{1}$ is smooth at the intersection points with the toric divisors.
[…]
Figure 1: Proof of Lemma 4.3, part (2), I
We perform the modification of the above tropicalization along the toric divisor ${\operatorname{Tor}}([(0,2),(2k,0)])$, in the sense of [20, Section 5] (see also [5, Section 2.2]). Recall that the curve $C$ has singularity of type $A_{2m}$ at $w$, and hence there exists a coordinate change
$x=x^{\prime},\quad y=y^{\prime}+(x^{\prime})^{k}+\sum_{k<j\leq m}c_{j}(x^{\prime})^{j}\ ,$ (16)
which converts the polynomial $F(x,y)$ into a polynomial
$F^{\prime}(x^{\prime},y^{\prime})=(y^{\prime})^{2}+a_{2m+1,0}(x^{\prime})^{2m+1}+\text{h.o.t.},\quad a_{2m+1,0}\neq 0\ .$
There exists a deformation of the coordinate change (16)
$x=x^{\prime},\quad y=y^{\prime}+(1+O(t_{1}))x^{k}+\sum_{k<j<m}(c_{j}+O(t_{1}))x^{j}\ ,$ (17)
which turns the family of polynomials $F_{t_{1}}(x,y)$, $|t_{1}|<{\varepsilon}$, into the family
$F^{\prime}_{t_{1}}(x^{\prime},y^{\prime})=\sum_{j=0}^{k-1}t_{1}^{(k-j)a}(a_{j1}+O(t_{1}))(x^{\prime})^{j}y^{\prime}+\sum_{j=2k-1}^{2m}t_{1}^{\nu(j)}(a_{j0}+O(t_{1}))(x^{\prime})^{j}$
[…]
The tropical limit of the latter family of polynomials defines the subdivision of the Newton polygon $P(F\_{t_{1}})$ into the triangles
$T_{1}={\operatorname{conv}}\{((0,2),(0,1),(2k-1,0)\},\quad T_{3}={\operatorname{conv}}\{(0,2),(2k-1,0),(2m+1,0)\}\ ,$
and the Newton polygon $P(F^{\prime})$ (see Figure 1(b)). Moreover, in principle, the triangle $T_{3}$ can be subdivided further. Taking into account that all limit curves associated with the considered tropical limit must be rational and that the coefficients of $F^{\prime}_{t_{1}}$ at the interior integral points of $T_{3}$ vanish, the triangle $T_{3}$ can be subdivided only into triangles of type ${\operatorname{conv}}\{(0,2),(2j^{\prime}-1,0),(2j^{\prime\prime}+1,0)\}$ with some $k\leq j^{\prime}\leq j^{\prime\prime}\leq m$. The function $\nu:\{2k-1,2k,...,2m+1\}\to{\mathbb{Z}}$ in the exponents of $t_{1}$ in (18) (with the extra value $\nu(2m+1)=0$) extends to a convex piecewise-linear function on the whole segment $[(2k-1,0),(2m+1,0)]$ with linearity domains induced by the subdivision of the triangle $T_{3}$. We make two remarks.
  • Since the germ at $\xi$ of the family {¯$\{\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<{\varepsilon}}$ is smooth and regularly parameterized by $t_{1}$ (see the beginning of Step (2)), we derive that $\mu(2m)=1$.
  • Since the linearity domains of $\nu$ are segments of even length, the value $\nu(2k-1)$ is even. On the other hand, formulas (17) yield that $\nu(2k-1)=\alpha$, and hence all exponents of $t_{1}$ in the leading terms of the coefficients of $F_{t_{1}}$ at $x^{j}y$, $10\leq j\leq 2k-$ 1, are even (see (14)). This means, in particular, that the limit curve $C_{1}$ given by (15), contributes to the curves ${\Phi}_{t}({\mathbb{P}}^{1})$ the same collection of real nodes with the same location with respect to ${\operatorname{Tor}}(\sigma)$ both for $t<t^{*}$ and for $t>t^{*}$. Hence, these singularities contribute the same factor to $W_{0}(\xi_{t})$ for $t<t^{*}$ and for $t>t^{*}$ (cf. formula (8)).
Now we analyze the contribution of the limit curves associated with the fragment $T_{3}$, and we consider the only case when $T_{3}$ is not subdivided. The other cases can be treated similarly. So, we have the limit curve
$C_{3}=\left\{(y^{\prime})^{2}+\sum_{j=2k-1}^{2m+1}a_{j0}(x^{\prime})^{j}=0\right\}\ ,$
which must be rational, and this yields that
$\sum_{j=2k-1}^{2m+1}a_{j0}(x^{\prime})^{j}=a_{2m+1,0}(x^{\prime})^{2k-1}Q(x^{\prime})^{2},\quad\deg Q=r:=m-k+1\ ,$
while the nodes of $C_{3}$ are $(\lambda_{j},0)$, $j=1,...,r$, with the numbers $\lambda_{j}$ being the roots of $Q$. Denote by $r_{+}$ (respectively, $r_{-}$) the number of positive (respectively, negative) roots of $Q$. Without loss of generality we can assume that $a_{2m+1,0}<0$. Then, the curve $C_{3}$ has $r_{+}$ hyperbolic nodes and $r_{-}$ elliptic nodes. In view of formulas (17), the nodes of $C_{3}$ yield the nodes
$(t_{1}(\lambda_{j}+O(t_{1})),\ t_{1}^{k}(\lambda_{j}^{k}+O(t_{1}))),\quad j=1,...,r\ ,$
of the curve ${\Phi}_{t}({\mathbb{P}}^{1})$, where $t_{1}=t-t^{*}$: for $t_{1}\lambda_{j}>0$ a hyperbolic node, for $t_{1}\lambda_{j}<0$ an elliptic one.
Thus, if $k$ is odd, then, for $t>t^{*}$ (respectively, $t<t^{*}$), the curve ${\Phi}_{t}({\mathbb{P}}^{1})$ gets $r_{+}$ (respectively, $r_{-}$) hyperbolic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ and $r_{-}$ (respectively, $r_{+}$) elliptic nodes outside ${\operatorname{Tor}}(P_{\Delta})^{+}$. None of these nodes contributes to the right-hand side of (8), and hence $W_{0}(\xi_{t})$ remains constant as $0<|t-t^{*}|<{\varepsilon}$.
If $k$ is even, then the curve ${\Phi}_{t}({\mathbb{P}}^{1})$ gets $r_{+}$ hyperbolic and $r_{-}$ elliptic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ if $t_{1}>0$, and gets $r_{-}$ hyperbolic and $r_{+}$ elliptic nodes in ${\operatorname{Tor}}(P_{\Delta})^{+}$ if $t_{1}<0$. Suppose that $r$ is even. Then $r_{+}\equiv r_{-}\mod 2$, and hence the first product in the right-hand side of formula (8) for $W_{0}(\xi_{t})$ remains constant. In turn, the orientation of the real local branch of ${\Phi}_{t}({\mathbb{P}}^{1})$ at $w$ remains the same (see Figure 2(a,b)), and hence the second product in the right-hand side of (8) remains constant as well. Suppose that $r$ is odd. Then $r_{+}\equiv r_{-}+1\mod 2$, and hence the first product in the right-hand side of formula (8) for $W_{0}(\xi_{t})$ changes sign as $t$ passes through $t^{*}$. However, the orientation of the real local branch of ${\Phi}_{t}({\mathbb{P}}^{1})$ at $w$ changes (see Figure 2(c,d)), which causes the change of sign in the second product in the right-hand side of (8).
[…]
Figure 2: Proof of Lemma 4.3, part (2), II
Thus, $W_{0}^{\kappa}(t)$ remains constant in the considered bifurcation.
@@
(3) Suppose that $\xi$ is as in Lemma 4.2(2ii). Here we choose a path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ defined in a neighborhood of $t^{*}$ so that the point $w_{i}^{\sigma}(t)$ is fixed and the point $w_{j}(t^{*})$ is mobile (together with some other point of $\widehat{\boldsymbol{w}}(t)$). Since $C$ is immersed, we extract the required local constancy of $W_{0}^{\kappa}(t)$ from the transversality relation (12) and the smoothness of the two following strata in ${\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$:
@@
(4) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)\in{\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}_{+}$. We closely follow the proof of [19, Theorem 5], providing here details for the reader’s convenience. Consider the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ locally obtained from $\widehat{\boldsymbol{w}}(t^{*})$ by moving one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{1}$ and one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{2}$, while fixing the position of the other points of $\widehat{\boldsymbol{w}}(t^{*})$. Then, we obtain a smooth one-dimensional germ in ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, which locally induces a deformation of the hyperbolic node ${\Phi}(\widehat{C},p)$ equivalent in suitable local coordinates in a neighborhood of $z={\Phi}(p)$ to $uv=\lambda(t)$ with $\lambda$ a smooth function in a neighborhood of $t^{*}$ such that $\lambda(t^{*})=0$ and $\lambda^{\prime}(t^{*})>0$ (see Figure 3(a)). The smoothness and the local deformation claims are straightforward from the (standard) considerations in the proof of [12, Lemma 11(2)], which reduce both claims to the following $h^{0}$-vanishing (cf. [12, Formula (16) and computations in the first paragraph in page 251]):
$h^{0}(\widehat{C}_{s},{\mathcal{O}}_{\widehat{C}_{s}}({\boldsymbol{d}}_{s}))=0,\quad s=1,2\ ,$ (19)
@@
$\deg{\boldsymbol{d}}_{s}=C_{s}^{2}-(C_{s}^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{s}]+2)-(c_{1}({\operatorname{Tor}}(P_{\Delta}))[C_{s}]-1)=-1>-2\ ,$
@@
Now we analyze the change of the refined invariants. For, we pass to the consideration of complex oriented curves. Denote by $\widehat{C}_{1}^{\pm}$ and $\widehat{C}_{2}^{\pm}$ the connected components of $\widehat{C}_{1}\setminus{\mathbb{R}}\widehat{C}_{1}$ and $\widehat{C}_{2}\setminus{\mathbb{R}}\widehat{C}_{2}$, respectively. When $t$ varies around $t^{*}$, the curve $\widehat{C}_{1}\cup\widehat{C}_{2}$ turns into ${\mathbb{P}}^{1}$, and we encounter two types of deformations: either $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{+},\widehat{C}_{2}^{-}$, respectively, or $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{-},\widehat{C}_{2}^{+}$, respectively. The type of the deformation agrees with the local complex orientation of the central curve as shown in Figure 3(b). This means that if for $t<t^{*}$ one encounters a deformation of the first type, then for $t>t^{*}$ it is of the second type, and vice versa. Set (cf. (8))
@@
Depending on the point of $C_{1}\cap C_{2}$ chosen for the smoothing, the invariant $(\kappa,{\varepsilon})$, where $\kappa$ is the quantum index and ${\varepsilon}$ is the second product in (8), changes as follows (see Figure 3(c)): either the real tangent vectors to $C_{1},C_{2}$ form a negative pair and
$\begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+}),&\end{cases}$
@@
$\begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-}).&\end{cases}$
For a simple geometric reason, the number of intersection points of the first kind equals the number of the intersection points of the second type. Hence, summing up over smoothings fo all real intersection points of $C_{1},C_{2}$, we derive the constancy of the numbers $W_{0}^{\kappa}(t)$ along the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ ($0<\zeta\ll 1$) in case of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ a birational immersion.

### v2

@@ 4.3 Wall-crossing
Figure 1: Proof of Lemma 4.3, part (2)
Since the deformation $\{C^{(t)}\}_{|t-t^{*}|<{\varepsilon}}$ is equigeneric, we can describe it via deformation of local parameterizations, see [31, Theorem 1 on page 73] and [3, Theorems 4.1 and 4.2]. Thus, $C\cap U=C^{(t^{*})}\cap U$ admits a (real) parametrization
$x=\tau^{2}+a_{2m+1}\tau^{2m+1}+\sum_{2\leq r\leq m}a_{2r}\tau^{2r}+\sum_{r\geq 2m-2k+2}a_{r}\tau^{r},\quad y=\tau^{2k},$
where $\tau\in({\mathbb{C}},0)$ and $a_{2m-2k+1}\neq 0$, and the curves $C^{(t)}\cap U$ are parameterized by
$x=\tau^{2}+a_{2m-2k+1}\tau^{2m-2k+1}+\sum_{2\leq r\leq m}a_{2r}\tau^{2r}+\sum_{r\geq 2m-2k+2}a_{r}\tau^{r}+\sum_{r\geq 1}b_{r}\tau^{r},\quad y=\tau^{2k},$
where $b_{r}=b_{r}(t)$ are analytic functions defined over ${\mathbb{R}}$ and vanishing at $t=t^{*}$.
The curves $C^{(t)}$, $t\neq t^{*}$, satisfy the conditions of Lemma 2.4, that is, $\alpha_{1}(t)>0$ for $0<|t-t^{*}|<{\varepsilon}$. Since $y=\tau^{2k}>0$ as $\tau\in{\mathbb{R}}^{*}$, all hyperbolic nodes are located in the domain $\{y>0\}$. Furthermore, a hyperbolic node corresponds to values $\tau_{1},\tau_{2}\in{\mathbb{R}}$, $\tau_{2}=-\tau_{1}$, such that $x(\tau_{)}=x(\tau_{2})$, which yields $x(\tau_{1})=\tau_{1}^{2}+O(\tau_{1}^{4})>0$. Consider an elliptic node of $C^{(t)}$, $t\neq t^{*}$. It corresponds to $\tau\in({\mathbb{C}},0)\setminus({\mathbb{R}},0)$ such tha $x(\tau),y(\tau)\in{\mathbb{R}}$. Let $k$ be odd. Suppose that $y(\tau)>0=\tau^{2k}>0$, i.e., $\tau=\theta\zeta_{2k}$, where $\theta>0$ and $(\zeta_{2k})^{2k}=1$, $\zeta_{2k}\neq\pm 1$. Then,
$x(\tau)=\alpha_{1}\theta\zeta_{2k}+(1+\alpha_{2})\theta^{2}(\zeta_{2k})^{2}+O(\theta^{3})\not\in{\mathbb{R}},$
since (recall that $k$ is odd)
$(1+\alpha_{2})\theta^{2}(\zeta_{2k})^{2}\not\in{\mathbb{R}}\quad\text{and}\quad\frac{\alpha_{1}\theta\zeta_{2k}}{(1+\alpha_{2})\theta^{2}(\zeta_{2k})^{2}}\not\in{\mathbb{R}}.$
Hence, all elliptic nodes live in the domain $\{y<0\}$. Let $k$ be even. Suppose that $y(\tau)=\tau^{2k}<0$, i.e., $\tau=\theta\eta\zeta_{2k}$, where $\theta>0$, $(\zeta_{2k})^{2k}=1$, $\eta=\exp\left(\frac{\pi\sqrt{-1}}{2k}\right)$. Thus,
$x(\tau)=\alpha_{1}\theta\eta\zeta_{2k}+(1+\alpha_{2})\theta^{2}\eta^{2}(\zeta_{2k})^{2}+O(\theta^{3}),$
and again we have
$(1+\alpha_{2})\theta^{2}(\eta^{2}\zeta_{2k})^{2}\not\in{\mathbb{R}}\quad\text{and}\quad\frac{\alpha_{1}\theta\zeta_{2k}}{(1+\alpha_{2})\theta^{2}\eta^{2}(\zeta_{2k})^{2}}\not\in{\mathbb{R}}.$
Hence, all elliptic nodes live in the domain $\{y>0\}$.
@@
(3) Suppose that $\xi$ is as in Lemma 4.2(2ii). Here we choose a path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ defined in a neighborhood of $t^{*}$ so that the point $w_{i}^{\sigma}(t)$ is fixed and the point $w_{j}(t^{*})$ is mobile (together with some other point of $\widehat{\boldsymbol{w}}(t)$). Since $C$ is immersed, we extract the required local constancy of $W_{0}^{\kappa}(t)$ and $\widetilde{W}_{0}^{\kappa}(t)$ from the transversality relation in the form of (15) which holds due to
$\deg{\boldsymbol{d}}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2)-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-2)=0>-2,$ (21)
and from the smoothness of the two following strata in ${\mathcal{O}}_{C,w}/{\mathfrak{m}}_{w}^{s}$:
@@
(4) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)\in{\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}_{+}$. We closely follow the proof of [24, Theorem 5], providing here details for the reader’s convenience. Consider the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ locally obtained from $\widehat{\boldsymbol{w}}(t^{*})$ by moving one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{1}$ and one point of $\widehat{\boldsymbol{w}}(t^{*})\cap C_{2}$, while fixing the position of the other points of $\widehat{\boldsymbol{w}}(t^{*})$. Then, we obtain a smooth one-dimensional germ in ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, which locally induces a deformation of the hyperbolic node ${\Phi}(\widehat{C},p)$ equivalent in suitable local coordinates in a neighborhood of $z={\Phi}(p)$ to $uv=\lambda(t)$ with $\lambda$ a smooth function in a neighborhood of $t^{*}$ such that $\lambda(t^{*})=0$ and $\lambda^{\prime}(t^{*})>0$ (see Figure 2(a)). The smoothness and the local deformation claims are straightforward from the (standard) considerations in the proof of [15, Lemma 11(2)], which reduce both claims to the following $h^{0}$-vanishing (cf. [15, Formula (16) and computations in the first paragraph in page 251]):
$h^{0}(\widehat{C}_{r},{\mathcal{O}}_{\widehat{C}_{r}}({\boldsymbol{d}}_{r}))=0,\quad r=1,2\ ,$ (22)
@@
$\deg{\boldsymbol{d}}_{r}=C_{r}^{2}-(C_{r}^{2}+C_{r}K_{{\operatorname{Tor}}(P_{\Delta})}+2)-(C_{r}K_{{\operatorname{Tor}}(P_{\Delta})}-1)=-1>-2\ ,$
@@
Now we analyze the change of the refined invariants, closely following the argument from [24, Section 6.4]. To this end, we pass to the consideration of complex oriented curves. Denote by $\widehat{C}_{1}^{\pm}$ and $\widehat{C}_{2}^{\pm}$ the connected components of $\widehat{C}_{1}\setminus{\mathbb{R}}\widehat{C}_{1}$ and $\widehat{C}_{2}\setminus{\mathbb{R}}\widehat{C}_{2}$, respectively. When $t$ varies around $t^{*}$, the curve $\widehat{C}_{1}\cup\widehat{C}_{2}$ turns into ${\mathbb{P}}^{1}$, and we encounter two types of deformations: either $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{+},\widehat{C}_{2}^{-}$, respectively, or $\widehat{C}_{1}^{+},\widehat{C}_{1}^{-}$ glue up with $\widehat{C}_{2}^{-},\widehat{C}_{2}^{+}$, respectively. The type of the deformation agrees with the local complex orientation of the central curve as shown in Figure 2(b). This means that if for $t<t^{*}$ one encounters a deformation of the first type, then for $t>t^{*}$ it is of the second type, and vice versa. Put (cf. (9) and (10))
@@
Depending on the point of $C_{1}\cap C_{2}$ chosen for the smoothing, the triple $(\kappa,{\varepsilon},\widetilde{\varepsilon})$, where $\kappa$ is the quantum index and ${\varepsilon},\widetilde{\varepsilon}$ denote the second product in (9), (10), respectively, changes as follows (see Figure 2(c)): either the real tangent vectors to $C_{1},C_{2}$ at the smoothed out intersection points form a negative pair and
$\begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+},\widetilde{\varepsilon}_{1}^{\;+}\widetilde{\varepsilon}_{2}^{\;+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-},\widetilde{\varepsilon}_{1}^{\;-}\widetilde{\varepsilon}_{2}^{\;-}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-},\widetilde{\varepsilon}_{1}^{\;+}\widetilde{\varepsilon}_{2}^{\;-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+},\widetilde{\varepsilon}_{1}^{\;-}\widetilde{\varepsilon}_{2}^{\;+}),&\end{cases}$
@@
$\begin{cases}(\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{-},\widetilde{\varepsilon}_{1}^{\;+}\widetilde{\varepsilon}_{2}^{\;-}),&\\ (-\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{+},\widetilde{\varepsilon}_{1}^{\;-}\widetilde{\varepsilon}_{2}^{\;+}),&\end{cases}\ \Longrightarrow\ \begin{cases}(\kappa_{1}+\kappa_{2},{\varepsilon}_{1}^{+}{\varepsilon}_{2}^{+},\widetilde{\varepsilon}_{1}^{\;+}\widetilde{\varepsilon}_{2}^{\;+}),&\\ (-\kappa_{1}-\kappa_{2},{\varepsilon}_{1}^{-}{\varepsilon}_{2}^{-},\widetilde{\varepsilon}_{1}^{\;-}\widetilde{\varepsilon}_{2}^{\;-}).&\end{cases}$
For a simple geometric reason, the number of intersection points of the first kind equals the number of the intersection points of the second type. Hence, summing up over smoothings fo all real intersection points of $C_{1},C_{2}$, we derive the constancy of the numbers $W_{0}^{\kappa}(t)$ and $\widetilde{W}_{0}^{\kappa}(t)$ along the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ ($0<\zeta\ll 1$) in case of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ a birational immersion.